# Mechanical detection of nuclear decays

### v1

@@ Abstract
More than a century of development has led to detectors that can precisely measure energy deposits from particles produced by radioactive decays. However, neutral particles emitted in such decays may escape detection if they do not subsequently interact within the detector. Here we report the detection of individual nuclear $\alpha$ decays through the mechanical recoil of the entire object in which the decaying nuclei are embedded. Momentum conservation ensures that such measurements are sensitive to any particles emitted in the decay, including neutral particles. Detection of the minuscule recoil of an object more than $10^{12}$ times more massive than the emitted particles is made possible by recently developed techniques in levitated optomechanics, which enable high-precision optical control and measurement of the mechanical motion of a micron-sized particle. The techniques developed here may find use in fields ranging from nuclear forensics to dark matter and neutrino physics.
@@ Introduction.
An alternative approach is to reconstruct the momentum imparted to the object containing the decaying nucleus by the recoil of its daughter, following the escape of the primary decay products [2]. For an object initially at rest, momentum conservation guarantees that the entire object must recoil with equal and opposite momentum to the total momentum carried by any escaping particles. This momentum-based reconstruction of particle decays has been demonstrated for individual recoiling nuclei [3, 4], but the tiny recoil has been previously undetectable for a much more massive solid object containing the decaying nucleus. For example, the nuclear recoil following an $\alpha$ decay will impart a momentum impulse of only $\sim 10^{-19}$ kg m/s to the object containing it, leading to a recoil of the micron-sized spheres studied here by only tens of nanometers.
@@
Detection of such tiny recoils requires extreme isolation of the object containing the decaying nucleus from thermal noise and precise measurement of the object’s motion. These stringent requirements can now be met by rapid progress in the optomechanical control and measurement of levitated particles [5]. By optically trapping nano- to micron-size particles in high vacuum, thermal noise can be made negligible [6]. The interaction of the trapped particle with the laser used to detect its position then provides the dominant noise source [7, 5, 6]. Ultimately, the minimal possible measurement induced noise is constrained by quantum mechanics [8, 9], and particles as large as a femtogram in mass (100 nm in diameter) are now reaching the quantum measurement regime [10, 11, 12, 13, 14, 15]. Such particles are extremely precise force sensors [16, 17], finding applications in tests of quantum mechanics with massive objects [18, 19, 20, 21, 22], searches for new short-distance interactions [23, 24], dark matter [25, 26, 27, 28], gravitational waves [29, 30], and neutrino physics [31]. Here we demonstrate the detection of individual nuclear $\alpha$ decays using optically levitated micron-sized spheres (with mass $>10$ pg). While these techniques already reach signal-to-noise ratios above 10, more than 100×100$\times$100 × further improvement in the momentum sensitivity is expected as these larger spheres are also brought into the quantum measurement regime.
@@
Figure 1: (a) ${}^{220}$n decay chain [32]. For each isotope, the decay half-life is given within the box. Arrows indicate the primary $\alpha$ decay energies and branching ratios (blue), as well as the relevant $\beta$ decay branching ratios (red). (b) Schematic of ${}^{212}$b implantation procedure described in the text. (c) Cross-sectional view of the electrodes surrounding the trap (an additional electrode on the front face and two weak imaging beams are not pictured). (d) Schematic of a sphere recoil following an $\alpha$ decay of ${}^{212}$i or ${}^{212}$o, for which the daughter nucleus is stopped within the sphere while the $\alpha$ (and additional low-energy $e^{-}$ and possibly secondary nuclear recoils, not pictured) exit the sphere. In this case, the momentum of the $\alpha$ particle can be inferred from the sphere recoil ($\vec{p}_{\alpha}=-\vec{p}_{sph}$).
@@
In our experiment, we detect single nuclear decays occurring in spherical silica particles of radius $\approx 1.5\ \mu$m that are optically trapped in vacuum at pressures between $10^{-8}$ to $10^{-7}$ mbar. Before trapping, the spheres are implanted with ${}^{212}$b, an unstable radon daughter in the thorium decay chain (Fig. 1 a). ${}^{220}$n is introduced into an implantation chamber [33] where it undergoes $\alpha$ decay to produce ${}^{216}$${}^{+}$ ions, which are drifted by an electric field to the sphere surfaces (Fig. 1 b). The $\alpha$ decay from the short-lived ${}^{216}$o ions implants the ${}^{212}$b daughters into the surfaces of the spheres at depths $\lesssim 60$ nm. The implanted spheres are then transferred to a high-vacuum chamber and loaded into an optical trap formed by a focused laser [33], where the subsequent decays are measured.
@@
The optical trap is surrounded by six planar electrodes (Fig. 1 c). The lower electrode is fixed to ground, while the upper five can be biased independently. The electrodes and trapping beam define a coordinate system with the trapping beam propagating vertically in the $z$ direction, and the $x$ and $y$ directions defined to be perpendicular to the electrode faces (Fig. 1 d). Once trapped in high vacuum, the sphere’s net electric charge is determined by monitoring the sphere’s motion in response to an oscillating electric field applied using these electrodes [7]. The net charge is calibrated into units of a single elementary charge, $e$, by observing discrete steps in the response [7, 34, 35], with the phase of the motion relative to the recorded electric field determining the polarity. Electrons can be added to the sphere via thermionic emission from a tungsten filament or removed from the sphere via an ultraviolet (UV) lamp. The trapping beam is circularly polarized to rotate the sphere to $\gtrsim 100$ kHz rotational frequencies, gyroscopically stabilizing its motion and eliminating noise from fluctuations in its angular orientation [7, 36]. The electrodes contain holes for the trapping beam and two weakly focused perpendicular imaging beams. The sphere position is detected by imaging the transmitted light on three detectors (one for each of $x$, $y$, and $z$[33]. The displacement sensitivity of this system is $\lesssim 10^{-11}$ m/$\sqrt{\mathrm{Hz}}$, which is sub-dominant to the force noise arising from residual fluctuations in the trapping beam pointing and intensity [7].
@@
Figure 2: (a, upper panel) Example of the measured charge versus time for a sphere implanted with ${}^{212}$b. Positive values of the charge correspond to a net excess of protons over electrons in the sphere. The black line shows the measured sphere charge (averaged in 12 s intervals) while the orange line shows the best fit to these data after reconstructing the location of each charge change. Light red vertical lines show when the filament adding $e^{-}$ to the sphere was active, while vertical gray bands indicate dead time during calibrations. The inset shows a zoom to several charge changes near the beginning of the data. (a, lower panel) Residual between the best-fit charge and the measured data. The residuals are typically $<1\ e$, demonstrating sensitivity to changes in the net charge of the sphere by a single $e$. (b) Example fit to the reconstructed number of charge changes versus time for the data shown on the left (blue points, with statistical error bars), with best fit $T_{1/2}=10.3^{+1.8}_{-1.5}$ h (blue band). (c) Distribution of reconstructed charge changes for all spheres studied in this work, which is empirically consistent with a sum (red line) of the two components (shaded) described in the text.
@@
The results presented here consist of data from six spheres implanted with ${}^{212}$b following the procedure above. Each sphere is continuously measured for 2–3 days after loading into the optical trap and pumping to high vacuum. Once in vacuum, the charge is continuously monitored and individual decays occurring within the sphere can be identified by detecting a change in the net charge of the sphere, $\Delta Q$. Figure 2 a shows an example of the net charge of the sphere over time after reaching $<10^{-7}$ mbar.
To minimize noise due to coupling of the sphere motion to stray electric fields, an automatic discharging procedure is implemented. When the net charge reaches $|Q|>50\ e$ either the tungsten filament or UV lamp is activated to reduce $|Q|$. Discharging periods are recorded by the data acquisition (DAQ) system and are excluded from the charge analysis (red lines in Fig. 2 a). Dead time due to the DAQ, impulse response calibrations, and these discharging periods is $<$10% of the total measurement time for all spheres considered.
For a typical implanted sphere, the decay frequency observed from these charge changes is initially around 2–5 decays per hour, with a decrease consistent with the ${}^{212}$b half-life, $T_{1/2}=10.6$ h (see Fig. 2 b). The background rate of charge changes was measured for a sphere that was not implanted with ${}^{212}$b. No charge changes were observed in 3 days of observation time following the initial neutralization of the sphere’s charge.
The distribution of observed charge changes from data across all spheres is shown in Fig. 2 c. While a detailed study of this distribution is beyond the scope of this work, several qualitative features are apparent. First, almost all charge changes are measured to be in the positive direction—i.e., the spheres nearly always lose more electrons than protons for both the $\alpha$ and $\beta$ decays in the decay chain. Second, two peaks in the charge distribution are observed, and the data are empirically consistent with a Gaussian-like component peaking around $\Delta Q=+2$ $e$ and a heavy-tailed component (here modeled by a log-normal distribution) peaking around $\Delta Q=+8$ $e$, but with a tail of events extending to $\Delta Q\gg+50$ e. Of the 257 total observed decays, 9 were observed with charge changes $>60\ e$ (beyond the range plotted in Fig. 2 c), with the largest observed change of $\Delta Q=+148\ e$.
@@
[…]
To accurately reconstruct the response of the sphere to such impulses, in situ calibrations of the impulse response for each sphere are performed in each direction using electric impulses of known amplitude. During these calibrations, the UV lamp and tungsten filament are first used to set the net charge of the sphere to $\pm$$e$, and repeated voltage pulses are applied sequentially to each of the 3 electrodes in the $x$, $y$, and $z$ directions. The pulse amplitude and duration are set to apply a total impulse of approximately 200 MeV/c, which is comparable to the expected impulse during an $\alpha$ decay. Using the precisely known charge and finite element method simulations of the electrode geometry, this calibration allows the impulse response to be measured to a relative accuracy of $<2$% in the $x$ and $y$ directions, and to $<10$% accuracy in the $z$ direction [33], limited by uncertainty in the knowledge of the sphere position relative to the electrodes.
@@
Several algorithms were studied to reconstruct the impulse amplitudes in the calibration data. An “optimal filter,” making use of the calibrated impulse response and measured noise spectrum for each sphere was found to provide the best resolution, although simpler reconstruction methods using only a bandpass filter or time domain fit provided signal-to-noise within a factor of 1.5$\times$ the optimal filter [33]. The resolution was found to be best for impulses applied in the $x$ direction, with a measured resolution in the range from $\sigma_{x}=15-27$ MeV/c for the spheres considered. This corresponds to an expected signal-to-noise ratio between $9-17$ for a ${}^{212}$o $\alpha$ decay with energy $E_{\alpha}=8.8$ MeV imparting its momentum $\sqrt{2m_{\alpha}E_{\alpha}}=256$ MeV/c along the $x$ direction, where $m_{\alpha}$ is the $\alpha$ particle mass. The resolution measured in the $y$ and $z$ directions is found to be poorer than in the $x$ direction in our current system due to the larger impact of technical noise (trapping laser pointing and amplitude fluctuations).
@@
Figure 3 shows an example candidate $\alpha$ decay coincident with an observed change in the net charge of the sphere. The impulse amplitude reconstructed from the optimal filter, and calibrated using the in situ calibration described above, shows a recoil of the sphere coincident in time with the charge change (whose time can be reconstructed with $<100$ ms precision), consistent with the impulse that would be expected for an $\alpha$ decay aligned between the $x$ and $y$ directions.
@@
The distribution of reconstructed impulse amplitudes in the $x$ direction coincident with detected charge changes for all spheres considered here is shown in Fig. 4. Compared to the charge data summarized in Fig. 2 c, this distribution contains fewer events (83 out of 257 total events) since only data periods at sufficiently low pressure that the sphere had reached a rotational velocity $>100$ kHz and the noise was stable were considered for the recoil analysis, while all time periods could be used in the charge analysis. Two of the six spheres were also not considered in this recoil analysis since they lacked in situ impulse calibration data. Only the $x$ projection of the momentum is considered in the recoil analysis due to its higher signal-to-noise (and observed systematic drifts in the $y$ amplitude between calibrations).
@@
Figure 4: Spectrum of the reconstructed $x$ impulse amplitudes for each observed charge change. All data are shown by the gray histogram, while events with $\Delta Q<30\ e$ are shown as blue points with error bars. The best fit to the spectrum (dark red line) is a sum of the predicted response from ${}^{212}$i $\alpha$ decays (dark orange shaded), ${}^{212}$o $\alpha$ decays (light orange shaded), and $\beta$ decays that impart momentum consistent with random sampling of the noise (gray shaded). The black dotted line shows the expected resolution for impulses in the $x$ direction measured from the in situ calibrations.
@@
In Fig. 4, all reconstructed recoils are shown by the gray line, while only recoils coincident with charge changes for which $\Delta Q<30\ e$ are indicated by the blue points. The reconstructed amplitudes are corrected for the small impulse induced by the charge monitoring field in the $x$ direction, which is estimated to provide negligible error after correction [33]. The data with $\Delta Q<30\ e$ are consistent with the expected distribution of momenta projected in the $x$ direction based on a Monte Carlo simulation of the $\alpha$ decays from ${}^{212}$i and ${}^{212}$o [33]. The maximum momentum transfer occurs when the $\alpha$ particle escapes along the $x$ direction and the daughter nucleus stops within the sphere.
@@
A small fraction of the observed decays, primarily with $\Delta Q>30\ e$, extend beyond the maximum momentum expected from these simulations. While further study is required to definitively identify the origin of these recoils, these events may arise from decays in which the nuclear recoil is emitted on a trajectory along the sphere’s surface, leading to a large number of secondary nuclear recoils ejecting Si and O atoms. These ejected atoms may carry some fraction of their momentum in the direction of the emitted $\alpha$, and might also be expected to produce the observed large changes in the net charge of the sphere as secondary electrons escape.
@@
Beyond fundamental physics, these techniques may find applications in nuclear forensics, which aims to determine the isotopic composition of a nuclear material [2, 39]. While future work is required to fully characterize what background levels may be possible, from the measurements presented here it appears plausible that sub-$\mu$Bq background rates will be achievable, possibly permitting detection of long-lived species such as ${}^{235}$ in single captured dust particles [2]. The measurements presented here will also allow the characterization of ejected daughter nuclei, as well as low-energy secondary electrons and ions by nuclear decays near a solid surface, which may be relevant in a variety of fields. For example, $\alpha$ therapy applications in nuclear medicine often aim to fully contain the subsequent nuclear recoils in micro- and nano-particles to avoid dispersal of toxic daughters in the body [40, 41]. Alternatively, radon daughters similar to the type studied here are commonly implanted in material surfaces by naturally occurring ${}^{222}$n, and their escape from surfaces (along with low energy secondaries) is relevant for a variety of rare event searches (e.g., [42, 43, 44]). Finally, the radon daughter decays ejecting $>100$ low-energy $e^{-}$ from a surface observed here may ultimately be observable in ion-based quantum computers, where a large number of charged particles passing through an ion array could simultaneously interact with multiple qubits [45]. Similar to phonon-mediated particle interactions in superconducting qubit arrays [46], these events may become evident only as error correction techniques begin to suppress uncorrelated errors and these systems are scaled to large size.
While this initial work has already demonstrated detection of single nuclear decays with signal-to-noise $>10$, substantial further improvement is expected. Future work is required to ensure the signal-to-noise is independent of the decay direction, characterize the background rates achievable, integrate conventional particle detectors around the trap, and to reach (and eventually surpass [47, 48]) the “standard quantum limit” for detection of the sphere recoil [9, 49]. The ongoing rapid progress in the field of levitated optomechanics promises to extend the future sensitivity of these techniques by orders-of-magnitude.

### v2

@@ Abstract
We report the detection of individual nuclear $\alpha$ decays through the mechanical recoil of the entire micron-sized particle in which the decaying nuclei are embedded. Momentum conservation ensures that such measurements are sensitive to any particles emitted in the decay, including neutral particles that may otherwise evade detection with existing techniques. Detection of the minuscule recoil of an object more than $10^{12}$ times more massive than the emitted particles is made possible by recently developed techniques in levitated optomechanics, which enable high-precision optical control and measurement of the mechanical motion of optically trapped particles. Observation of a change in the net charge of the particle coincident with the recoil allows decays to be identified with background levels at the micro-Becquerel level. The techniques developed here may find use in fields ranging from nuclear forensics to dark matter and neutrino physics.
@@ Introduction.
An alternative approach is to reconstruct the momentum imparted to the object containing the decaying nucleus by the recoil of its daughter, following the escape of the primary decay products [2]. This momentum-based reconstruction is sensitive to any escaping particles, including neutral particles, and has been demonstrated for individual recoiling nuclei [3, 4]. However, the tiny recoil has been previously undetectable for a much more massive solid object containing the decaying nucleus. For example, the $\alpha$ decays detected here impart a momentum impulse of only $\sim 10^{-19}$ kg m/s to the micron-sized spheres containing the decaying nuclei.
@@
Detection of such tiny recoils requires extreme isolation of the object containing the decaying nucleus from thermal noise and precise measurement of the object’s motion. These stringent requirements can now be met by rapid progress in the optomechanical control and measurement of levitated particles [5]. By optically trapping nano- to micron-size particles in high vacuum, thermal noise can be made negligible [6]. The interaction of the trapped particle with the laser used to detect its position then provides the dominant noise source [7, 5, 6]. Ultimately, the minimal possible measurement induced noise is constrained by quantum mechanics [8, 9], and particles as large as a femtogram in mass (100 nm in diameter) are now reaching the quantum measurement regime [10, 11, 12, 13, 14, 15]. Such particles are extremely precise force sensors [16, 17], finding applications in tests of quantum mechanics with massive objects [18, 19, 20, 21, 22], searches for new short-distance interactions [23, 24], dark matter [25, 26, 27, 28], gravitational waves [29, 30], and neutrino physics [31].
Here we demonstrate the detection of individual nuclear $\alpha$ decays using optically levitated micron-sized spheres (with mass $>10$ pg). Decays are detected through the change in the net electric charge of the particle following the decay and the coincident recoil of the entire particle. The combination of these signatures allows both extremely low backgrounds and sensitivity to the momentum carried by the decay products through the measurement of the recoil. While these techniques already reach signal-to-noise ratios above 10 for the recoil measurement, more than 100×100$\times$100 × further improvement in the momentum sensitivity is expected as these larger spheres are also brought into the quantum measurement regime.
@@
Figure 1: (a) 220Rn decay chain [32]. For each isotope, the decay half-life is given within the box. Arrows indicate the primary $\alpha$ decay energies and branching ratios (blue), as well as the relevant $\beta$ decay branching ratios (red). (b) Schematic of 212Pb implantation procedure described in the text. (c) Cross-sectional view of the electrodes surrounding the trap (an additional electrode on the front face and two weak imaging beams are not pictured). (d) Schematic of a sphere recoil following an $\alpha$ decay of 212Bi or 212Po, for which the daughter nucleus is stopped within the sphere while the $\alpha$ (and additional low-energy $e^{-}$ and possibly secondary nuclear recoils, not pictured) exit the sphere. In this case, the momentum of the $\alpha$ particle can be inferred from the sphere recoil ($\vec{p}_{\alpha}=-\vec{p}_{sph}$).
@@
In our experiment, we detect single nuclear decays occurring in silica spheres of radius $\approx 1.5\ \mu$m that are optically trapped in vacuum at pressures between $10^{-8}$ to $10^{-7}$ mbar. Before trapping, the spheres are implanted with 212Pb, an unstable radon daughter in the thorium decay chain (Fig. 1 a). 220Rn is introduced into an implantation chamber [33] where it undergoes $\alpha$ decay to produce 216Po+ ions, which are drifted by an electric field to the sphere surfaces (Fig. 1 b). The 216Po $\alpha$ decays then implant 212Pb daughters into the spheres at depths $\lesssim 60$ nm. The implanted spheres are then transferred to a high-vacuum chamber and loaded into an optical trap formed by a focused laser, where the subsequent decays are measured.
@@
The optical trap is surrounded by six planar electrodes (Fig. 1 c). The lower electrode is grounded, while the upper five can be biased independently. The electrodes and trapping beam define a coordinate system with the trapping beam propagating vertically in the $z$ direction, and the $x$ and $y$ directions defined to be perpendicular to the electrode faces (Fig. 1 d). Once trapped in high vacuum, the sphere’s net electric charge is determined by monitoring the sphere’s motion in response to an oscillating electric field applied using these electrodes [7]. The net charge is calibrated into units of a single elementary charge, $e$, by observing discrete steps in the response [7, 34, 35], with the phase of the motion relative to the electric field determining the polarity. Electrons can be added to the sphere via thermionic emission from a tungsten filament or removed from the sphere via an ultraviolet (UV) lamp. The trapping beam is circularly polarized to rotate the sphere to $\gtrsim 100$ kHz rotational frequencies, gyroscopically stabilizing its motion and eliminating noise from fluctuations in its angular orientation [7, 36]. The electrodes contain holes for the trapping beam and two weakly focused perpendicular imaging beams. The sphere position is detected by imaging the transmitted light on three detectors (one for each of $x$, $y$, and $z$). The dominant noise source in the $x$ direction at the base pressure of $2\times 10^{-8}$ mbar is thermal noise from residual gas. The $y$ and $z$ directions suffer from additional technical noise due to fluctuations in the trapping beam pointing and intensity [7].
@@
Figure 2: (a, upper panel) Example of the measured charge versus time for a sphere implanted with 212Pb. Positive values of the charge correspond to a net excess of protons over electrons in the sphere. The black line shows the measured sphere charge (averaged in 12 s intervals) while the orange line shows the best fit to these data after reconstructing the location of each charge change. Light red vertical lines show when the filament adding $e^{-}$ to the sphere was active, while vertical gray bands indicate dead time during calibrations. The inset shows a zoom to several charge changes near the beginning of the data. (a, lower panel) Residual between the best-fit charge and the measured data. The residuals are typically $<1\ e$, demonstrating sensitivity to changes in the net charge of the sphere by a single $e$. (b) Example fit to the reconstructed number of charge changes versus time for the data shown on the left (blue points, with statistical error bars), with best fit $T_{1/2}=10.3^{+1.8}_{-1.5}$ h (blue band). (c) Distribution of reconstructed charge changes for all spheres studied in this work, which is empirically consistent with a sum (red line) of the two components (shaded) described in the text.
@@
The results presented here consist of data from six spheres implanted with 212Pb following the procedure above. Each sphere is continuously measured for 2–3 days after loading into the optical trap and pumping to high vacuum. Once in vacuum, the charge is continuously monitored and individual decays occurring within the sphere are identified by detecting a change in the net charge of the sphere, $\Delta Q$. Figure 2 a shows an example of the net charge of the sphere over time after reaching $<10^{-7}$ mbar.
An automatic discharging procedure is implemented to maintain the net charge $|Q|<50\ e$. Discharging periods are recorded by the data acquisition (DAQ) system and excluded from the charge analysis (red lines in Fig. 2 a). Dead time due to the DAQ, impulse response calibrations, and these discharging periods is $<$10% of the total measurement time for all spheres considered.
For a typical implanted sphere, the decay frequency observed from these charge changes is initially around 2–5 decays per hour, with a decrease consistent with the 212Pb half-life, $T_{1/2}=10.6$ h (see Fig. 2 b). The background rate of charge changes was measured for an unimplanted sphere, with no charge changes observed in 3 days. As a result, any charge changes can be used to identify decays with backgrounds at the $<1/(\mathrm{day})\sim\mu$Bq level. For the measurements presented here, this charge measurement is crucial to reject transient noise bursts [33], which were found in previous work to be associated with vibrational or acoustic noise [27].
The distribution of charge changes from data across all spheres is shown in Fig. 2 c. While a detailed study of this distribution is beyond the scope of this work, several qualitative features are apparent. First, almost all charge changes are measured to be in the positive direction—i.e., the spheres nearly always lose more electrons than protons for both the $\alpha$ and $\beta$ decays in the decay chain. Second, two peaks in the charge distribution are observed, and the data are empirically consistent with a Gaussian-like component peaking around $\Delta Q=+2$ $e$ and a heavy-tailed component (here modeled by a log-normal distribution) peaking around $\Delta Q=+8$ $e$, but with a tail of events extending to $\Delta Q\gg+50$ e. Of the 257 total observed decays, 9 were observed with charge changes $>60\ e$ (beyond the range plotted in Fig. 2 c), with the largest observed change of $\Delta Q=+148\ e$.
@@
The observed charge changes are used to tag decay events and provide timing information to search for recoils from the momentum transferred to the sphere during a decay. Since the time for all decay products to exit or stop within the sphere is much faster than its $\sim$10 ms mechanical response time, the decay will impart an essentially instantaneous momentum impulse, which can be detected through the subsequent motion of the sphere’s center-of-mass.
To reconstruct the response of the sphere to such impulses, in situ calibrations are performed using electric impulses of known amplitude. Using the known net charge of the sphere and finite element method simulations of the electrode geometry, this calibration allows the impulse response to be measured to a relative accuracy of $<2$% in the $x$ and $y$ directions, and to $<10$% accuracy in the $z$ direction, limited by uncertainty in the knowledge of the sphere position relative to the electrodes [33].
@@
An “optimal filter” [37, 33], making use of the calibrated impulse response and measured noise spectrum for each sphere, was found to provide the best resolution for impulse amplitude reconstruction. In the $x$ direction, a resolution in the range of $\sigma_{x}=15-27$ MeV/c was measured for the spheres considered. This corresponds to an expected signal-to-noise ratio between $9-17$ for a 212Po $\alpha$ decay with energy $E_{\alpha}=8.8$ MeV imparting its momentum $\sqrt{2m_{\alpha}E_{\alpha}}=256$ MeV/c along the $x$ direction, where $m_{\alpha}$ is the $\alpha$ particle mass. The resolution measured in the $y$ and $z$ directions is found to be poorer than in the $x$ direction in our current system due to the larger impact of technical noise.
@@
Figure 3 shows a candidate $\alpha$ decay coincident with an observed change in the charge of the sphere. The impulse amplitude reconstructed from the optimal filter, and calibrated using the in situ calibration described above, shows a recoil of the sphere coincident in time with the charge change (whose time can be reconstructed with $<100$ ms precision), consistent with the impulse that would be expected for an $\alpha$ decay aligned between the $x$ and $y$ directions.
@@
The distribution of reconstructed impulse amplitudes in the $x$ direction coincident with detected charge changes for all spheres considered here is shown in Fig. 4. Compared to the charge data summarized in Fig. 2 c, this distribution contains fewer events (83 out of 257 total events) since only data periods at sufficiently low pressure that the sphere had reached a rotational velocity $>100$ kHz and the noise was stable were considered for the recoil analysis, while all time periods were used in the charge analysis. Two of the six spheres were also not considered in this recoil analysis since they lacked in situ impulse calibrations. Only the $x$ projection of the momentum is considered in the recoil analysis due to its higher signal-to-noise (and observed systematic drifts in the $y$ amplitude between calibrations).
@@
Figure 4: Spectrum of the reconstructed $x$ impulse amplitudes for each observed charge change. All data are shown by the gray histogram, while events with $\Delta Q<30\ e$ are shown as blue points with error bars. The best fit to the spectrum (dark red line) is a sum of the predicted response from 212Bi $\alpha$ decays (dark orange shaded), 212Po $\alpha$ decays (light orange shaded), and $\beta$ decays that impart momentum consistent with random sampling of the noise (gray shaded). The black dotted line shows the expected resolution for impulses in the $x$ direction measured from the in situ calibrations.
@@
In Fig. 4, all reconstructed recoils are shown by the gray line, while only recoils coincident with charge changes for which $\Delta Q<30\ e$ are indicated by the blue points. The reconstructed amplitudes are corrected for the small impulse induced by the charge monitoring field in the $x$ direction, which is estimated to provide negligible error after correction. The data with $\Delta Q<30\ e$ are consistent with the expected distribution of momenta projected in the $x$ direction based on a Monte Carlo simulation of the $\alpha$ decays from 212Bi and 212Po [33]. In addition to these $\alpha$ decays, the decay chain contains several $\beta$ and $\gamma$ transitions, which may occur in coincidence with the $\alpha$ decay. While these particles can carry non-negligible energy compared to the $\alpha$, their relative momentum is much lower ($\sim$1–3 MeV/c), and is not detectable with the resolution of the existing setup. The maximum momentum transfer occurs when the $\alpha$ particle escapes along the $x$ direction and the daughter nucleus stops within the sphere.
@@
A small fraction of the observed decays extend beyond the maximum momentum expected from these simulations. The cut removing events with $\Delta Q<30\ e$ in Fig. 4 indicates that only events in the tail of the charge change distribution contribute to these unexpectedly large decays. Systematic checks were performed to ensure these large reconstructed recoils were not due to miscalibration or reconstruction errors. While further study is required to definitively identify the origin of these recoils, these events may arise from decays in which the nuclear recoil is emitted along the sphere’s surface, leading to a large number of secondary nuclear recoils ejecting Si and O atoms. These ejected atoms may carry some fraction of their momentum in the direction of the emitted $\alpha$, and might also produce the observed large changes in the net charge of the sphere as secondary electrons escape.
@@
Beyond fundamental physics, these techniques may find applications in nuclear forensics, which aims to determine the isotopic composition of a nuclear material [2, 40]. While future work is required to fully characterize what background levels may be possible, from the measurements presented here it appears plausible that sub-$\mu$Bq background rates will be achievable, possibly permitting detection of long-lived species such as 235U in single captured dust particles [2]. The measurements presented here will also allow the characterization of ejected radon daughters, as well as low-energy secondary electrons and ions, from decays near a solid surface, which may be relevant for applications in nuclear medicine [41, 42] and rare event searches (e.g., [43, 44, 45]). Finally, the radon daughter decays ejecting $>100$ low-energy $e^{-}$ from a surface observed here may ultimately be observable in ion-based quantum computers, where a large number of charged particles passing through an ion array could simultaneously interact with multiple qubits [46]. Similar to phonon-mediated particle interactions in superconducting qubit arrays [47], these events may become evident only as error correction techniques begin to suppress uncorrelated errors and these systems are scaled to large size.
While this initial work has already demonstrated detection of single nuclear decays with signal-to-noise $>10$, substantial further improvement is expected. Future work is required to ensure the signal-to-noise is independent of the decay direction, characterize the background rates achievable, integrate conventional particle detectors around the trap, and to reach—and eventually surpass [48, 49]—the “standard quantum limit” for the detection of the sphere recoil [50, 9], where the momentum resolution would be $\sqrt{\hbar m\omega_{0}}\approx 50$ keV/c for the spheres considered here (with mass $m\sim 10$ pg and angular resonant frequency $\omega_{0}\sim 2\pi\times 100$ Hz). The ongoing rapid progress in the field of levitated optomechanics promises to extend the future sensitivity of these techniques by orders-of-magnitude.